# BeDKD: Backdoor Defense based on Dynamic Knowledge Distillation and Directional Mapping Modulator

### v1

@@ Abstract
Although existing backdoor defenses have gained success in mitigating backdoor attacks, they still face substantial challenges. In particular, most of them rely on large amounts of clean data to weaken the backdoor mapping but generally struggle with residual trigger effects, resulting in persistently high attack success rates (ASR). Therefore, in this paper, we propose a novel Backdoor defense method based on Directional mapping module and adversarial Knowledge Distillation (BeDKD), which balances the trade-off between defense effectiveness and model performance using a small amount of clean and poisoned data. We first introduce a directional mapping module to identify poisoned data, which destroys clean mapping while keeping backdoor mapping on a small set of flipped clean data. Then, the adversarial knowledge distillation is designed to reinforce clean mapping and suppress backdoor mapping through a cycle iteration mechanism between trust and punish distillations using clean and identified poisoned data. We conduct experiments to mitigate mainstream attacks on three datasets, and experimental results demonstrate that BeDKD surpasses the state-of-the-art defenses and reduces the ASR by 98 without significantly reducing the CACC. Our code are available in https://github.com/CAU-ISS-Lab/Backdoor-Attack-Defense-LLMs/tree/main/BeDKD.
1 Introduction
@@
In recent years, deep neural networks (DNNs) have achieved great success in the field of natural language processing (NLP), such as sentiment analysis [1, 2], machine translation [3, 4] and natural language generation [5, 6]. However, recent studies show that DNNs are highly vulnerable to backdoor attacks [7, 8, 9, 10].
@@
Backdoor attacks generally introduce an invisible vulnerability in DNNs, allowing attackers to control or manipulate the model’s output when the input contains the specific trigger patterns [11, 12]. To carry out a backdoor attack, the attacker first injects triggers into a small amount of clean data to poison the training set, and then trains the victim model. In inference, the poisoned model responds normally to clean data, while it responds incorrectly to poisoned data based on the attacker’s target label. The prevalence of backdoor attacks poses significant security risks to deep neural networks [13, 14, 15, 16].
@@
To defend against backdoor attacks, researchers have explored many backdoor defense methods, broadly categorized into data-level [17, 18, 19, 20] and model-level [21, 22, 23] approaches. As shown in Figure 1(a) and (b), the goal of data-level methods is to identify poisoned data, while the goal of model-level methods is to erase the backdoor of the poisoned model. The former identifies poisoned data from the input data via external models or fine-tuned models. Even though these methods have achieved success in mitigating backdoor attacks, their primary strategy is to avoid activating backdoors rather than essentially eliminate backdoors. In contrast, the later mainly erases backdoors through data cleaning, training, knowledge distillation (KD), or neuronal pruning. Although the existing model-level methods remove backdoors effectively, they reduce the accuracy of the poisoned model on the clean data. Therefore, achieving a satisfactory trade-off between backdoor defense and maintaining model performance remains a significant challenge.
@@
More recently, some defense methods have been introduced to alleviate the above trade-off problem. Zhao et al. [24] randomly flips the label of a clean proxy dataset to fine-tune the poisoned model, enabling it to identify poisoned data. To erase backdoors, Zhao et al. [45] leverages a clean proxy dataset to fine-tune the BERT and uses the fine-tuned BERT as the teacher model, which guides the poisoned student model to unlearn the backdoors via knowledge distillation. Although they excel at both mitigating backdoor attacks and preserving model performance, they require quantities of clean data to fine-tune models, limiting their application in the real world.
@@
We explore a novel model-level backdoor defense method based on directional mapping module and adversarial knowledge distillation (BeDKD), which makes a satisfied trade-off between defense effectiveness and model performance via a small amount of clean and poisoned data.
@@
2 Related Work
2.1 Backdoor Attack
Dai et al. [26] and Chen et al. [27] insert meaningful fixed short sentences and rare words into clean data. To improve the stealthiness of triggers, Qi et al. [28] and Pan et al. [29] rewrite sentences with a specific syntactic structure and style. Yan et al. [BITE] capitalize on spurious correlations between the target label and specific words in training data. To further improve stealthiness and text quality, Du et al. [30] fine-tune LLMs based on attribute control to generate poisoned data. Similarly, Li et al. [31] design hand-crafted prompt to guide LLMs to generate rephrased poisoned data. With the advancement of backdoor attacks, designing an accurate and effective backdoor defense is still a critical and pressing challenge.
2.2 Backdoor Defense
(1) Data-Level Defenses. Qi et al. [32] utilize an external language model as a grammar outlier detector to remove trigger words from the input. Yang et al. [33] use an additional prompt-based optimizer to verify the output logit permutation. Chen et al. [17] identify trigger words using word importance scores. Due to the poisoned model’s sensitivity to triggers, Gao et al. [18] detect poisoned data by randomly perturbing features and analyzing output changes of each data. Similarly, He et al. [34] used gradients or self-attention scores to self-defend against backdoor attacks. Although existing data-level defenses successfully defend against backdoor attacks, they still have live backdoors. (2) Model-Level Defenses. He et al. [35] compute the spurious correlation between text features and labels to clean the poisoned training set and retain the victim model. Zhao et al. [36] erase backdoors through attention head pruning and weight- normalization. Pei et al. [23] train multiple classifiers on divided sub-training sets and ensemble their predictions. These defenses mitigate backdoor attacks effectively, while they struggle to balance the defense trade-off and require substantial clean data for fine-tuning.
2.3 Knowledge Distillation
Knowledge distillation (KD) compresses larger or ensemble networks (teacher models) into smaller networks (student models) [37]. Feature maps and attention mechanisms have proven effective in KD, enabling student models to learn high-quality intermediate representations from teacher models, thereby enhancing distillation and improving performance [38, 39]. KD has been applied to speech recognition [40, 41], visual recognition [42, 43], backdoor defense [44, 45]. Zhao et al. [45] fine-tune BERT on a large task-related clean dataset as the teacher model to guide the poisoned model to erase backdoors via knowledge distillation. However, they rely heavily on large volumes of clean data, posing challenges in low-resource scenarios.
3 Methodology
3.1 Preliminaries
@@
Following the previous backdoor defenses [17, 23, 45], the defender is user. The defender has access to the training set but is unaware of the presence of poisoned data within it. The goal of defender is to distill a clean model using the downloaded poisoned dataset, while preserving the clean mapping and eliminating the backdoor mapping. This means that the defended model should have a low attack success rate on the poisoned test set, while maintaining a high classification accuracy on the clean test set.
3.2 Overview of BeDKD
@@
3.3 Distilled DMM for Locating Poisoned Data
Traditional backdoor defenses use clean data for fine-tuning or distillation to erase the backdoors [36, 45]. However, they require a large number of clean data and fall short of completely eliminating the backdoor mapping (higher ASR). This paper leverages a small number of clean samples to identify a small number of poisoned samples and incorporates them into the distillation process, enabling the model to more effectively remove backdoors. To find poisoned samples, we propose the Directional Mapping Module (DMM), which has the same structure as the poisoned model and is distilled by a small amount of flipped clean data to disrupt the clean mapping of the DMM while reinforcing the backdoor mapping, thereby facilitating the identification of trustworthy poisoned samples. The goal of DMM is to make the probability distribution difference of clean mapping as large as possible, while making the probability distribution difference of backdoor mapping as small as possible.
Assume that we have access to a small number of clean data [50, 36, 45]. We modify the ground-truth label of clean data and flip it to an incorrect label to create a flipped clean data , where is label space. We initialized the DMM with shared parameters from the .
@@ 3.3 Distilled DMM for Locating Poisoned Data
3.4 Adversarial Knowledge Distillation
Traditional knowledge distillation focuses on guiding the student model to learn the feature distributions of the teacher model, thereby facilitating knowledge transfer and enhancing generalization [46]. However, in the task of backdoor defense, directly applying traditional knowledge distillation methods can lead the student model to simultaneously learn both the clean mapping and backdoor mapping from the poisoned teacher model, making it difficult to eliminate backdoors (detailed discussion in Section 4.3). In addition, although some studies utilize task-related clean datasets to distill a clean model from the poisoned model, such as W2SDefense [45], they require a large amount of clean data, which limits their practical application. To address this issue, we propose an Adversarial Knowledge Distillation (AKD) method, which employs an adversarial distillation strategy to promote the learning of clean mapping while suppressing the learning of backdoor mapping on limited clean and poisoned data (as shown in Figure 2(c)). Specifically, the teacher model is the poisoned model with frozen parameters, while the student model () shares the same architecture and parameters as . The AKD framework adopts a cycle iteration mechanism, performing trust distillation on a small amount of clean data and punish distillation on a small amount of poisoned data identified in the previous step. By alternating between the two types of distillation, the backdoor mapping is eliminated without reducing the clean mapping.
@@ 3.4 Adversarial Knowledge Distillation
The algorithm of the BeDKD is listed in Appendix B. During the training stage, the AKD performs a cycle iteration mechanism, alternating between trust and punish distillation. By alternating these two distillations, AKD ensures that the clean mapping is strengthened through the trust distillation, while the backdoor mapping is gradually erased during the punish distillation.
4 Evaluation
4.1 Evaluation Settings
Datasets Attacks.
Baselines Metrics.
We compare BeDKD with five mainstream defenses: Fine-Tuning (FT) [50], ONION [32], IMBERT [34], TextGuard [23], and W2SDefense [45]. Details are listed in Appendix E. To be fair, we follow previous studies and utilize four commonly adopted metrics. ASR measures the accuracy of poisoned models on poisoned data. CACC assesses the accuracy of both poisoned and clean models on clean data. FAR represents the percentage of poisoned data classified as clean out of all poisoned data. FRR indicates the percentage of clean data classified as poisoned out of all clean data.
@@ 4.1 Evaluation Settings
We leverage the AdamW optimizer with the learning rate of 3 to train the poisoned model (widely used BERT) for 10 epochs. According to previous experience, the temperatures of the DMM and AKD are set to 1.5 and 2.5, respectively. The and are both set to 0.3. We train the DMM and AKD for 20 epochs and 50 epochs. More details are shown in Appendix F.
4.2 Comparison Results
Table 1 summarizes the performance comparison of BeDKD with baselines. "No Defense" means the poisoned models without any defenses. All backdoor attacks always achieve more than 99 ASR. The proposed BeDKD significantly outperforms all baselines on most attack settings and lowers around 98 of all backdoor attacks without compromising CACC in most cases. For insertion-based attacks, BadWords and AddSent leverage visible rare words and fixed sentences as triggers, respectively. Although most baselines can mitigate these attacks, BeDKD achieves lower ASR and higher CACC, especially the average ASR and CACC on the three datasets achieve 0.01 and 88.41, which is better than the best baseline, W2SDefense (average ASR 15.92 and CACC 87.83). For paraphrase-based attacks, SynBkd, StyBkd, AttrBkd, and BGMAttack utilize invisible syntax templates, style, attribution, and AI-generated texts as trigger patterns, respectively. BeDKD still surpasses the best baselines and reduces the average ASR to 1.93 (16.14 than W2SDefense). These results show that BeDKD effectively defends against both visible and invisible trigger patterns. On the OLID dataset, all defense baselines cannot work well because the scale of the dataset is small. While BeDKD still effectively defends against all backdoor attacks on the OLID dataset and reduces the average ASR to 1.27 (17.61 than W2SDefense). Overall, BeDKD makes a satisfactory trade-off on a small amount of clean data. More victim experiments are listed in Appendix G.
4.3 Ablation Study
@@ 4.4 Sensitivity Analysis
As shown in Table 6, BeDKD can effectively defend against backdoor attacks through clean proxy data (HateSpeech) and AI-generated texts (GPT-4o). For clean HateSpeech and AI-generated texts, the average ACCs are 82.72 and 77.95, which are close to the CACC of OLID 82.12. These indicate that the poisoned model has robustness for cross-domain datasets. After defensive, the average ASRs are reduced to 1.02 (HateSpeech) and 1.54 (AI-generated Texts). Compared with HateSpeech, the CACC of AI-generated texts is lower at 79.42. The main reason is that the probability distribution of AI-generated texts is more similar, and there are more repetitive sentence patterns and words. These results indicate that BeDKD has strong generalization and robustness. BeDKD does not rely on same domain data and can still effectively mitigate backdoor on cross-domain data.
5 Conclusion

### v2

# BeDKD: Backdoor Defense Based on Directional Mapping Module and Adversarial Knowledge Distillation

@@ Abstract
Although existing backdoor defenses have gained success in mitigating backdoor attacks, they still face substantial challenges. In particular, most of them rely on large amounts of clean data to weaken the backdoor mapping but generally struggle with residual trigger effects, resulting in persistently high attack success rates (ASR). Therefore, in this paper, we propose a novel Backdoor defense method based on Directional mapping module and adversarial Knowledge Distillation (BeDKD), which balances the trade-off between defense effectiveness and model performance using a small amount of clean and poisoned data. We first introduce a directional mapping module to identify poisoned data, which destroys clean mapping while keeping backdoor mapping on a small set of flipped clean data. Then, the adversarial knowledge distillation is designed to reinforce clean mapping and suppress backdoor mapping through a cycle iteration mechanism between trust and punish distillations using clean and identified poisoned data. We conduct experiments to mitigate mainstream attacks on three datasets, and experimental results demonstrate that BeDKD surpasses the state-of-the-art defenses and reduces the ASR by 98 without significantly reducing the CACC.
Code — https://github.com/CAU-ISS-Lab/Backdoor-Attack-Defense-LLMs/tree/main/BeDKD
Introduction
@@
In recent years, deep neural networks (DNNs) have achieved great success in the field of natural language processing (NLP), such as sentiment analysis (Wang et al. 2020; Huang et al. 2023), machine translation (Wang et al. 2021, 2024) and natural language generation (Sun et al. 2023; Vice et al. 2024). However, recent studies show that DNNs are highly vulnerable to backdoor attacks (Li et al. 2022a, b; Wan et al. 2024; Nguyen et al. 2024).
@@
Backdoor attacks generally introduce an invisible vulnerability in DNNs, allowing attackers to control or manipulate the model’s output when the input contains the specific trigger patterns (He et al. 2022; Wu et al. 2022). To carry out a backdoor attack, the attacker first injects triggers into a small amount of clean data to poison the training set, and then trains the victim model. In inference, the poisoned model responds normally to clean data, while it responds incorrectly to poisoned data based on the attacker’s target label. The prevalence of backdoor attacks poses significant security risks to deep neural networks (Rahman et al. 2020; Ma et al. 2021; Tiwari, Yuan, and Zhang 2022; Zhu et al. 2022).
@@
To defend against backdoor attacks, researchers have explored many backdoor defense methods, broadly categorized into data-level (Chen and Dai 2021; Gao et al. 2022; Xi et al. 2023; Li et al. 2023) and model-level (Jin, Wang, and Shang 2022; Zhao, Xu, and Yuan 2024; Pei et al. 2024) approaches. As shown in Figure 1(a) and (b), the goal of data-level methods is to identify poisoned data, while the goal of model-level methods is to erase the backdoor of the poisoned model. The former identifies poisoned data from the input data via external models or fine-tuned models. Even though these methods have achieved success in mitigating backdoor attacks, their primary strategy is to avoid activating backdoors rather than essentially eliminate backdoors. In contrast, the later mainly erases backdoors through data cleaning, training, knowledge distillation (KD), or neuronal pruning. Although the existing model-level methods remove backdoors effectively, they reduce the accuracy of the poisoned model on the clean data. Therefore, achieving a satisfactory trade-off between backdoor defense and maintaining model performance remains a significant challenge.
@@
More recently, some defense methods have been introduced to alleviate the above trade-off problem. Zhao et al. (Zhao et al. 2024a) randomly flips the label of a clean proxy dataset to fine-tune the poisoned model, enabling it to identify poisoned data. To erase backdoors, Zhao et al. (Zhao et al. 2024b) leverages a clean proxy dataset to fine-tune the BERT and uses the fine-tuned BERT as the teacher model, which guides the poisoned student model to unlearn the backdoors via knowledge distillation. Although they excel at both mitigating backdoor attacks and preserving model performance, they require quantities of clean data to fine-tune models, limiting their application in the real world.
@@
We explore a novel model-level backdoor defense based on directional mapping module and adversarial knowledge distillation (BeDKD), which makes a satisfied trade-off between defense effectiveness and model performance via a small amount of clean and poisoned data.
@@
Related Work
Backdoor Attack
Dai et al. (Dai, Chen, and Li 2019) and Chen et al. (Chen et al. 2021) insert meaningful fixed short sentences and rare words into clean data. To improve the stealthiness of triggers, Qi et al. (Qi et al. 2021b) and Pan et al. (Pan et al. 2022) rewrite sentences with a specific syntactic structure and style. Yan et al. (Yan, Gupta, and Ren 2022) capitalize on spurious correlations between the target label and specific words in training data. To further improve stealthiness and text quality, Du et al. (Du et al. 2024) fine-tune LLMs based on attribute control to generate poisoned data. Similarly, Li et al. (Li et al. 2024) design hand-crafted prompt to guide LLMs to generate rephrased poisoned data. With the advancement of backdoor attacks, designing an accurate and effective backdoor defense is still a critical challenge.
Backdoor Defense
(1) Data-Level Defenses. Qi et al. (Qi et al. 2021a) utilize an external language model as a grammar outlier detector to remove trigger words from the input. Yang et al. (Yang et al. 2021) use an additional prompt-based optimizer to verify the output logit permutation. Chen et al. (Chen and Dai 2021) identify trigger words using word importance scores. Due to the poisoned model’s sensitivity to triggers, Gao et al. (Gao et al. 2022) detect poisoned data by randomly perturbing features and analyzing output changes of each data. Similarly, He et al. (He et al. 2023a) used gradients or self-attention scores to self-defend against backdoor attacks. Although existing data-level defenses successfully defend against backdoor attacks, they still have live backdoors. (2) Model-Level Defenses. He et al. (He et al. 2023b) compute the spurious correlation between text features and labels to clean the poisoned training set and retain the victim model. Zhao et al. (Zhao, Xu, and Yuan 2024) erase backdoors through attention head pruning and weight- normalization. Pei et al. (Pei et al. 2024) train multiple classifiers on divided sub-training sets and ensemble their predictions. These defenses mitigate backdoor attacks effectively, while they struggle to balance the defense trade-off and require substantial clean data for fine-tuning.
Knowledge Distillation
Knowledge distillation (KD) compresses larger or ensemble networks (teacher models) into smaller networks (student models) (Hinton, Vinyals, and Dean 2015). Feature maps and attention mechanisms have proven effective in KD, enabling student models to learn high-quality intermediate representations from teacher models, thereby enhancing distillation and improving performance (Byeongho Heo 2019; Tian, Krishnan, and Isola 2020). KD has been applied to speech recognition (Zhao et al. 2020; Zhang, Liu, and Liu 2023), visual recognition (Zagoruyko and Komodakis 2017; Zhao and Han 2021), backdoor defense (Li et al. 2021; Zhao et al. 2024b). Zhao et al. (Zhao et al. 2024b) fine-tune BERT on a large task-related clean dataset as the teacher model to guide the poisoned model to erase backdoors via knowledge distillation. However, they rely heavily on large volumes of clean data, posing challenges in low-resource scenarios.
Methodology
Preliminaries
@@
Following the previous backdoor defenses (Chen and Dai 2021; Pei et al. 2024; Zhao et al. 2024b), the defender is user. The defender has access to the training set but is unaware of the presence of poisoned data within it. The goal of defender is to distill a clean model using the downloaded poisoned dataset, while preserving the clean mapping and eliminating the backdoor mapping. This means that the defended model should have a low attack success rate on the poisoned test set, while maintaining a high classification accuracy on the clean test set.
Overview of BeDKD
@@
Distilled DMM for Locating Poisoned Data
Traditional backdoor defenses use clean data for fine-tuning or distillation to erase the backdoors (Zhao, Xu, and Yuan 2024; Zhao et al. 2024b). However, they require a large number of clean data and fall short of completely eliminating the backdoor mapping (higher ASR). This paper leverages a small number of clean samples to identify a small number of poisoned samples and incorporates them into the distillation process, enabling the model to more effectively remove backdoors. To find poisoned samples, we propose the Directional Mapping Module (DMM), which has the same structure as the poisoned model and is distilled by a small amount of flipped clean data to disrupt the clean mapping of the DMM while reinforcing the backdoor mapping, thereby facilitating the identification of trustworthy poisoned samples. The goal of DMM is to make the probability distribution difference of clean mapping as large as possible, while making the probability distribution difference of backdoor mapping as small as possible.
Assume that we have access to a small number of clean data (Yosinski et al. 2014; Zhao, Xu, and Yuan 2024; Zhao et al. 2024b). We modify the ground-truth label of clean data and flip it to an incorrect label to create a flipped clean data , where is label space. We initialized the DMM with shared parameters from the .
@@
Adversarial Knowledge Distillation
Traditional knowledge distillation focuses on guiding the student model to learn the feature distributions of the teacher model, thereby facilitating knowledge transfer and enhancing generalization (Phuong and Lampert 2019). However, in backdoor defense task, directly applying traditional knowledge distillation can lead the student model to simultaneously learn both the clean and backdoor mapping from the poisoned teacher model, making it difficult to eliminate backdoors (detailed discussion in Section Ablation Study). In addition, although some studies utilize task-related clean datasets to distill a clean model from the poisoned model, such as W2SDefense (Zhao et al. 2024b), they require a large amount of clean data, which limits their practical application. To address this issue, we propose an Adversarial Knowledge Distillation (AKD), which employs an adversarial distillation strategy to promote the learning of clean mapping while suppressing backdoor mapping on limited clean and poisoned data (as shown in Figure 2(c)). Specifically, the teacher model is the poisoned model with frozen parameters, while the student model () shares the same architecture and parameters as . The AKD adopts a cycle iteration mechanism, performing trust distillation on a small amount of clean data and punish distillation on a small amount of poisoned data identified in the previous step. By alternating between two types of distillation, the backdoor mapping is eliminated without reducing the clean mapping.
@@
During the training stage, the AKD performs a cycle iteration mechanism, alternating between trust and punish distillation. By alternating these two distillations, AKD ensures that the clean mapping is strengthened through the trust distillation, while the backdoor mapping is gradually erased during the punish distillation.
Evaluation
Evaluation Settings
Datasets & Attacks.
We conduct experiments on SST2 (Socher et al. 2013), AGnews (Zhang, Zhao, and LeCun 2015), and OLID (Dai et al. 2020). We simulate six prominent backdoor attacks: AddSent (Dai, Chen, and Li 2019), BadWords (Chen et al. 2021), SynBkd (Qi et al. 2021b), StyBkd (Pan et al. 2022), AttrBkd (Du et al. 2024), and BGMAttack (Li et al. 2024).
Baselines & Metrics.
We compare BeDKD with five mainstream defenses: Fine-Tuning (FT) (Yosinski et al. 2014), ONION (Qi et al. 2021a), IMBERT (He et al. 2023a), TextGuard (Pei et al. 2024), and W2SDefense (Zhao et al. 2024b).To be fair, we follow previous studies and utilize four commonly adopted metrics. ASR and CACC measure the accuracy of poisoned data and clean data. FAR means the percentage of poisoned data classified as clean out of all poisoned data. FRR means the percentage of clean data classified as poisoned out of all clean data.
@@
We leverage the AdamW optimizer with the learning rate of 3 to train the poisoned model (widely used BERT) for 10 epochs. According to previous experience, the temperatures of the DMM and AKD are set to 1.5 and 2.5, respectively. The and are both set to 0.3. We train the DMM and AKD for 20 epochs and 50 epochs.
Comparison Results
Table 1 summarizes the performance comparison of BeDKD with baselines. ”No Defense” means the poisoned models without any defenses. All backdoor attacks always achieve more than 99 ASR. BeDKD significantly outperforms baselines on most attack settings and lowers around 98 of all backdoor attacks without compromising CACC in most cases. For insertion-based attacks, BadWords and AddSent use visible rare words and fixed sentences as triggers. Although most baselines can mitigate them, BeDKD achieves lower ASR and higher CACC, especially the average ASR and CACC on three datasets achieve 0.01 and 88.41, which is better than the best baseline, W2SDefense (average ASR 15.92 and CACC 87.83). For paraphrase-based attacks, SynBkd, StyBkd, AttrBkd, and BGMAttack use invisible syntax templates, style, attribution, and AI-generated texts as triggers. BeDKD still reduces the average ASR to 1.93 (16.14 than W2SDefense). These results show that BeDKD effectively mitigates both visible and invisible triggers. On the OLID dataset, all defense baselines cannot work well because the small scale of the dataset. While BeDKD still effectively defends against all backdoor attacks on OLID dataset and reduces the average ASR to 1.27 (17.61 than W2SDefense). Moreover, BeDKD requires 232s to achieve defense on SST2, which is dramatically more efficient, running orders of magnitude faster than W2SDefense (4,000s) and TextGuard (30,000s). Overall, BeDKD makes a satisfactory trade-off on a small amount of clean data.
Ablation Study
@@
As shown in Table 6, BeDKD can effectively defend against backdoor attacks through clean proxy data (HateSpeech) and AI-generated texts (GPT-4o). For clean HateSpeech and AI-generated texts, the average ACCs are 82.72 and 77.95, which are close to the CACC of OLID 82.12. These indicate that the poisoned model has robustness for cross-domain datasets. After defensive, the average ASRs are reduced to 1.02 (HateSpeech) and 1.54 (AI-generated Texts). Compared with HateSpeech, the CACC of AI-generated texts is lower at 79.42. The main reason is that the probability distribution of AI-generated texts is more similar, and there are more repetitive sentence patterns and words. These results show that BeDKD has strong generalization and robustness.
Conclusion